Define the solutions to
Q(y)=y^4-10y^2+9
a. (y-1)(y_1)(y-3)(y+3)
b. 1, -1, 3, -3
c. -1, -3 only
d. None of the above

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

Q(y)=y^4-10y^2+9
Q(Y)=(Y^2-9)(Y^2-1)
Q(Y)=(Y+3)(Y-3)(Y+1)(Y-1)
Y+3=0 OR Y=-3
Y-3=0 OR Y=3
Y+1=0 OR Y=-1
Y-0=0 OR Y=1
Answer b) is the correct answer.

Answer by stanbon(75887)

b
Cheers,
Stan H.